Given the linear equation y= -5/6+2, find the y-coordinates of the points (-6), (0), and (6). Plot those points and graph the linear equation.
Answer by jim_thompson5910(28595)

y= -5/6x+2

y= -5/6(-6)+2

y= 30/6+2

y= 5+2

y= 7

So we have one point (-6,7)

---------------
y= -5/6x+2

y= -5/6(0)+2

y= 0/6+2

y= 0+2

y= 2

So we have one point (0,1)
---------------
y= -5/6x+2

y= -5/6(6)+2

y= -30/6+2

y= -5+2

y= -3

So we have one point (6,-3)


Now plot these points and draw a straight line through them to graph the line (I'll let you do this).